finance

QUESTION 3
a) On 1 January 2017, Gyamera Limited granted 100 share options to each of its 300 employees, with each of the share options being conditional upon the employee working for Gyamera Limited until 31 December 2019. At the grant date, the FV of each share option was GHC15.00. During 2017, 15 employees left Gyamera Limited and the company’s directors estimated that a total of 20% of the 300 employees would leave during the three-year period 2017-2019. At the beginning of 2018, Gyamera Limited modified the terms and conditions of the share option by reducing the exercise price. This had the effect of increasing the FV of a share option at the beginning of 2018 by GHC9.00.
During 2018, a further six employees left the company and the directors revised their estimate of the total number of the 300 employees to 15% that would leave the company during the three-year period 2017-2019.
During 2019, a further five employees left the company.
Required;
Calculate the remuneration expense that should be recognised in Gyamera Limited’s financial statements in respect of the share-based payment agreement for each of the three years 2017,2018 and 2019.                  (6 Marks)

b) IAS 32 makes it clear that the following items are not financial instruments.

i. Physical assets, eg inventories, property, plant and equipment, leased assets and intangible assets (patents, trademarks etc)


Page 7 of 7

ii. Prepaid expenses, deferred revenue and most warranty obligations iii. Liabilities or assets that are not contractual in nature (Deferred revenue, warranty obligations

Required
Write a Memo to your CEO, Mr Francis Tagoe explaining with reasons why the items listed above do not qualify as financial instruments.         (6 Marks)

c) Your managing director, Mr Nicholas Adoboe-Mensah is not convinced why your company should give something back to the society, Explain with examples to him in a form of a memo the links between corporate performance, and corporate social responsibility.    (5 Marks)


d) Global Reporting Initiatives structures key performance indicators according to a hierarchy category, aspect and indicator. Indicators are grouped in terms of the three dimensions of the conventional definition of sustainability – economic, environment, and social.
Required
Write a memo to your Chief Executive officer Mr Baba Ahmed explaining to him the meaning of Global Reporting Initiatives and state two examples each of the aspect of each of the dimensions
Answer #1

Financial Year - 2017

Total Estimated cost = 300*100*15 = 450000

Total number of estimated employee leaving the firm = 20% of strength

So Total Estimated expense = 80% of 450000 = 360000

Expenses to be recordd in 2017 = 360000/3 = 120000


Financial Year -2018

Total Estimated Cost = 300*100*9 = 270000

Total number of estimated employee leaving the firm = 15% of strength

So Total Estimated Espense = 270000* 80 % = 216000

Expenses to be recorded in 2018 = 216000/3 = 72000

Expenditure over booked in 2017 = 120000 - 72000 = 48000

Net Expenses to be booked in 2018 = 72000 - 48000 = 24000


Financial Year - 2019

Total Estimated Cost = 270000

Actual Number of Employees left the firm = 26

So Total Expnses = 274*100*9=246600

Expenses booked till date = 120000 + 24000 = 144000

Expenses to br booked in 2019 = 246600-144000 = 10600


answered by: whoami